A rectangle is 6 times as long as it is wide. The area is 54 square feet. What is the perimeter of the rectangle?

Answer by checkley79(3341)

L=6W
A=LW
54=6W*W
54=6W^2
W^2=54/6
W^2=9
W=SQRT9
W=3 ANS. FOR THE WIDTH.
L=6*3=16 ANS. FOR THE LENGTH.
P=2L+2W
P=2*16+2*3
P=32+6=42 ANS. FOR THE PERIMETER.